OpenStudy (anonymous):

Find three consecutive even integers whose sum is 114.

OpenStudy (anonymous):

x,x+1,x+2 x+(x+1)+(x+2)=144 3x+3=144 3x=144/3=48 x=48/3=16

OpenStudy (anonymous):

so x=16

OpenStudy (anonymous):

Isn't it 3 consecutive integers?

OpenStudy (anonymous):

yes. x 1st , x+1 2nd , x+3 3rd

OpenStudy (anonymous):

so , it is 16, 16+1, 16+3? 16, 17, 19

OpenStudy (anonymous):

no 16,16+1,16+2

OpenStudy (anonymous):

that is 16,17&18

OpenStudy (anonymous):

But their sum does not equal 114

OpenStudy (anonymous):

sorry that is x+2 3rd

OpenStudy (anonymous):

Everytime I use x+2 on my third, I get 16, 17, 19 and they aren't consecutive and their sum doesn't add to 114

OpenStudy (anonymous):

sorry i have some mistake in calculation

OpenStudy (anonymous):

x+(x+1)+(x+2)=144 3x+3=144 3x=144-3=141 x=141/3=47

OpenStudy (anonymous):

37, 38 ,39

OpenStudy (anonymous):

sum has to be 114 not 144 though

OpenStudy (anonymous):

3x+3=114 3x=114-3=111 x=111/3=37

OpenStudy (anonymous):

now cleared

OpenStudy (anonymous):

YES! I HAD IT RIGHT!!!

OpenStudy (anonymous):

i have many careless mistake saorry for that

OpenStudy (anonymous):

It's alright, I make them too